A brass rod is 2m at a certain temperature. What is it's length for a temperature rise of 100k, if the coefficient of linear expansion of brass is 8×10^-5 k^-1?


1
Expert's answer
2021-05-10T10:07:03-0400

The change in length with temperature occurs according to the following equation:


"L=L_0+\Delta L,\\
\Delta L=L_0\alpha\Delta T,\\
L=L_0(1+\alpha\Delta T),\\L=2(1+8·10^{-5}·100)=2.016\text{ m}."